Mechanics

A uniform rod of length l, hinged at the lower end is free to rotate in a vertical plane. If the rod is held vertically in the beginning and then released, the angular acceleration of the rod when it makes an angle of 45°with the horizontal is ?

Biswayan Mukherjee
7 Years agoGrade

1 Answer

Arun
7 Years ago
 

When it makes 45 degree with the horizontal then at this position torque acting on rod is =mgl/2 cos45

                                                                                                                             =mgl/2√2

We know that

Torque=moment of inertia * angular acceleration

mgl/2√2  =ml2/3  * angular acceleration         where m is the mass of rod

angular acceleration =3g/2√2 l